OpenStudy (anonymous):

a gift box has a square base with a length of 30 cm. What is the approximate lenght of the diagonal of the base of the gift box

OpenStudy (kittiwitti1):

Find the hypotenuse using the Pythagorean Theorem\[a^2 + b^2 = c^2\]\[c = \sqrt{a^2+b^2} = sqrt{30^2 + 30^2} \approx 42 cm\]